OpenStudy (anonymous):

Use rules of logarithms to expan. Log (x+3)^2/ (x-2)(x^2+5)^4.

OpenStudy (anonymous):

\[ \log \frac{(x+3)^2}{ (x-2)(x^2+5)^4.}\] \[ \log \frac{(x+3)^2}{ (x-2)(x^2+5)^4.} = \log (x+3)^2 - \log[ (x-2)(x^2+5)^4]\] \[ =2 \log (x+3) - [\log (x-2)+ \log (x^2+5)^4]\] \[=2 \log (x+3) - [\log (x-2)+ 4 \log (x^2+5)]\] \[=2 \log (x+3) - \log (x-2)- 4 \log (x^2+5)\] is the expanded form of the given problem. @spontaneous44

OpenStudy (anonymous):

@spontaneous44

OpenStudy (anonymous):

thank you so much I understand it better!

OpenStudy (anonymous):

@spontaneous44 Medal please!!!

OpenStudy (anonymous):

I'm new here how do i give medals?

OpenStudy (ikram002p):

@spontaneous44 i think it solved now :))

OpenStudy (anonymous):

yes :)